# Physically Valid Biomolecular Interaction Modeling with Gauss-Seidel Projection

### v1

@@ Abstract
Biomolecular interaction modeling has been substantially advanced by foundation models, yet they often produce all-atom structures that violate basic steric feasibility. We address this limitation by enforcing physical validity as a strict constraint during both training and inference with a unified module. At its core is a differentiable projection that maps the provisional atom coordinates from the diffusion model to the nearest physically valid configuration. This projection is achieved using a Gauss-Seidel scheme, which exploits the locality and sparsity of the constraints to ensure stable and fast convergence at scale. By implicit differentiation to obtain gradients, our module integrates seamlessly into existing frameworks for end-to-end finetuning. With our Gauss-Seidel projection module in place, two denoising steps are sufficient to produce biomolecular complexes that are both physically valid and structurally accurate. Across six benchmarks, our $2$-step model achieves the same structural accuracy as state-of-the-art $200$-step diffusion baselines, delivering ${\sim}10\times$ wall-clock speedups while guaranteeing physical validity.
@@ 1 Introduction
End-to-end, all-atom protein structure predictors that integrate deep learning with generative modeling are emerging as transformative tools for biomolecular interaction modeling (Senior et al., 2020b; Jumper et al., 2021; Abramson et al., 2024; Corso et al., 2023; Watson et al., 2023). These systems achieve unprecedented accuracy in predicting arbitrary biomolecular complexes and are shaping the future of computational biology and drug discovery.
Contrary to their high structural accuracy, current predictors often fail to satisfy a conceptually simpler but basic requirement: the physical validity of the all-atom output. This failure, also commonly termed hallucination, appears as steric clashes, distorted covalent geometry, and stereochemical errors (Fig. 8). In practice, however, physical validity is a prerequisite: Atom-level physics violations hinder expert assessment (Senior et al., 2020a), undermine structure-based reasoning and experimental planning (Lyu et al., 2019), and also destabilize downstream computational analyses such as molecular dynamics (Hollingsworth & Dror, 2018; Lindorff-Larsen et al., 2011).
The root cause of this problem lies in the design of current predictors: as generative models, they are trained to match the empirical distribution of known structures, without enforcing physical validity as a strict constraint in the training objective. Consequently, these models can assign non-zero probability to non-physical configurations, a flaw that persist even with large-scale training, both in data and model size, as observed in recent works (Wohlwend et al., 2024; Passaro et al., 2025; Buttenschoen et al., 2024; Team et al., 2025). Methods such as Boltz-1-steering (Wohlwend et al., 2024) reintroduce physics as inference-time guidance by biasing the sampling process towards physically valid regions. Such guidance can reduce violations, but still cannot guarantee validity: with finite guidance strength and update steps, invalid configurations remain reachable.
@@
Concretely, we instantiate this decoupling with a differentiable Gauss-Seidel projection placed after the denoiser. It solves a constrained optimization that projects all-atom coordinates to the nearest physically valid configuration. The constraints are tightly coupled through shared atoms and their number scales with the number of atoms, making the optimization problem large. Within the training loop, the module must also converge in a handful of iterations with low memory per forward and backward pass. These requirements make pure first-order gradient descent methods impractical, which requires tiny steps and many iterations to converge. We therefore adopt a Gauss-Seidel scheme (Saad, 2003), which sweeps over all constraints for a few iterations, enforcing each locally by updating only the affected atoms. It leverages locality and sparsity of the constraints and thereby yields faster, more stable convergence than gradient descent. The projection module is differentiable via implicit differentiation. We integrate it as a drop-in layer within existing biomolecular interaction frameworks (e.g., Boltz (Wohlwend et al., 2024)) and finetune end-to-end. At inference time, 2 denoising steps suffice to produce structurally accurate, physically valid protein complexes.
We evaluate our method on six protein-complex benchmarks: CASP15, Test (Wohlwend et al., 2024), PoseBusters (Buttenschoen et al., 2024), AF3-AB (Abramson et al., 2024), dsDNA, and RNA-Protein (Ma et al., 2025). Comparisons include $200$-step generative models (Boltz-1, Boltz-1-Steering (Wohlwend et al., 2024), Boltz-2 Passaro et al. (2025), and Protenix (Team et al., 2025)) and the few-step baseline Protenix-Mini (Gong et al., 2025). Despite using only two denoising steps, our model achieves competitive structural accuracy while guaranteeing physical validity. For runtime, our model delivers a ${\sim}10\times$ wall-clock speedup over baselines. Our study largely closes the gap between guaranteed physical validity and state-of-the-art structural accuracy under few-step sampling, enabling 2-step all-atom predictions with an order-of-magnitude faster inference.
@@ 2 Related Work
Protein Structure Prediction. Deep learning-based foundation models are reshaping the task of protein structure prediction that produces 3D coordinates from input sequences (Senior et al., 2020b; Jumper et al., 2021). End-to-end variants extended this paradigm to complexes and large-scale training and inference (Baek et al., 2021; Evans et al., 2022). AlphaFold-3 further moved to all-atom biomolecular interaction modeling with a diffusion sampler (Abramson et al., 2024), followed by improvements on generative models, such as Boltz-2 (Passaro et al., 2025) and Protenix (Team et al., 2025). These methods typically require hundreds of denoising steps. Few-step variants such as Protenix-Mini (Team et al., 2025) reduce sampling to two steps while maintaining accuracy, yet still lack guarantees on physical validity. Boltz-1-Steering (Wohlwend et al., 2024) is the first to explicitly target this issue by steering sampling with physics-informed potentials, but the guidance is soft and cannot preclude violations. By contrast, we introduce a Gauss-Seidel projection layer that enforces physical validity for both training and inference.
@@
Diffusion Models with Tilted Distribution Sampling. Diffusion models have demonstrated strong capabilities in producing samples given input conditions (Ho et al., 2020; Dhariwal & Nichol, 2021; Song et al., 2021). However, in many cases, users require the generated samples to satisfy certain specific constraints. There are primarily two types of constraint-aware diffusion methods: (i) Feynman-Kac (FK) steering-based conditional sampling methods (Trippe et al., 2023; Singhal et al., 2025) , which steer the sampling toward regions that meet desired conditions through reweighting of the path distribution in the diffusion process, and (ii) manifold-based diffusion models (Bortoli et al., 2022; Elhag et al., 2024) , which explicitly construct a constrained manifold, ensuring that generated samples inherently satisfy the desired constraints. However, FK-steering typically requires a large number of sampling steps, and explicit manifold construction can be computationally expensive when constraints are complex. We introduce a Gauss-Seidel projection module that enforces constraints in training and inference, thereby circumventing both excessive sampling iterations and the need for expensive manifold construction.
@@
Constraint Enforcement via Position-Based Dynamics. Enforcing physical constraints has long been central to physics simulation across computational physics and computer graphics. To handle boundary conditions (Macklin et al., 2014), collisions, contact, friction (Bridson et al., 2002), and bond stretching (Ryckaert et al., 1977), many methods have been developed to keep simulations physically consistent. Among these, position-based dynamics and its extensions (Müller et al., 2007; Macklin & Müller, 2016; Chen et al., 2024) iteratively project particle positions to satisfy predefined constraints, proving effective for soft bodies, fluids, and coupled phenomena (Bender et al., 2014). By updating positions directly instead of integrating stiff forces, these methods remain stable under large timesteps and avoid costly global solves. Since constraints are enforced through small, local projections, the overall complexity scales nearly linearly with the number of constraints, which aligns with our goal of generating proteins that respect local physical constraints.
@@ 3 Preliminary
All-Atom Diffusion Model. Proteins can be represented at all-atom resolution (Abramson et al., 2024; Wohlwend et al., 2024; Passaro et al., 2025), where a structure is specified by the Cartesian coordinates of all atoms $\hat{\mathbf{x}}\in\mathbb{R}^{N\times 3}$, with $N$ the total number of atoms. This representation enables direct modeling of side-chain orientations, local packing, and explicit inter-atomic geometry. AlphaFold3 (Abramson et al., 2024) introduced an atom-level diffusion approach that iteratively denoises atomic coordinates from Gaussian noise. Conditioned on features produced by the atom-attention encoder, MSA module, and PairFormer, a denoising network runs for hundreds of steps (typically 200) to generate the all-atom prediction $\hat{\mathbf{x}}$.
Physics-Guided Steering. All-atom diffusion sampling is stochastic and therefore cannot guarantee physical validity. To address this, Boltz-1-Steering (Wohlwend et al., 2024), built upon the Feynman-Kac framework (Singhal et al., 2025), employs physics-informed potentials. At each denoising step, these potentials score the validity and tilt the learned denoising model toward a lower-energy configuration. The energy is a weighted sum of seven terms: tetrahedral atom chirality, bond stereochemistry, planarity of double bonds, internal ligand distance bounds, steric clashes avoidance, non-overlapping symmetric chains, and the preservation of covalently bonded chains. Since this tilted distribution cannot be sampled directly, the method uses importance sampling to draw multiple candidates, which are then refined with gradient descent updates to further reduce violations. Although this steering reduces physical errors, it remains a soft approach: it biases the sampling process toward validity during inference rather than embedding validity into the training.
@@ 4.1 Penalty-based Formulation for Physical Validity
where $\mathbf{C}(\cdot)\!:\!\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}^{m}$ concatenates all constraints that characterize physical validity, and its zero set defines the feasible space. We follow Boltz-1-Steering (Wohlwend et al., 2024) to define the constraints. Detailed formulation is provided in Appendix G. Since all terms are locally defined on pairs or four-atom groups, the total number of constraints $m$ scales with atoms and can be large. For instance, for PDB 8TGH which consists of $470$ residues with $7,082$ atoms, $m$ is ${\sim}13$M.
A standard approach to solving the constrained problem Eq. 1 is the penalty method (Boyd & Vandenberghe, 2004), in which the constraints are incorporated as an exterior penalty into the objective:
@@
where we choose a quadratic penalty $E(\cdot)$ and $\bm{\alpha}\in\mathbb{R}^{m\times m}$ is a block-diagonal matrix of penalty coefficients. For sufficiently small $\bm{\alpha}$ (i.e., large weights $\bm{\alpha}^{-1}$), Eq. 2 serves as a numerical realization of the hard-constrained projection: the minimizer satisfies the constraint up to numerical tolerance. In our implementation, we set $\bm{\alpha}_{j}=10^{-7}$ to enforce a tight feasibility threshold.
@@ 4.2 Gauss-Seidel Constraint Projection
Solving Eq. 2 alone is classical and admits many off-the-shelf solvers. Embedding it as a differentiable module within a biomolecular modeling pipeline, however, imposes stricter requirements: very fast convergence per forward pass and training step, and stable convergence under stiff, tightly coupled constraints. Pure first-order gradient descent, as used in Boltz-1-Steering (Wohlwend et al., 2024), is ill-suited: it requires small step sizes and prohibitively many iterations to converge to a zero constraint penalty, making unrolled end-to-end training slow. We therefore propose a Gauss-Seidel projection that exploits locality and sparsity by sequentially updating only the atoms affected by each constraint, thereby achieving fast and stable convergence.
@@
We then introduce the Lagrange multiplier $\bm{\lambda}(\mathbf{x})\!:=\!-\bm{\alpha}^{-1}\mathbf{C}(\mathbf{x})$ by following Stuart & Humphries (1996) and Servin et al. (2006) to obtain the coupled system:
@@
with updates $\mathbf{x}^{(n+1)}=\mathbf{x}^{(n)}+\Delta\mathbf{x}$ and $\bm{\lambda}^{(n+1)}=\bm{\lambda}^{(n)}+\Delta\bm{\lambda}$. It can be shown that with initialization $\mathbf{x}^{(0)}=\hat{\mathbf{x}}$ and $\bm{\lambda}^{(0)}=\mathbf{0}$, the linearized iterates converge to a solution of the original nonlinear system Eq. 4. See Appendix D for a proof. Applying the Schur complement (Zhang, 2005) with respect to $\mathbf{I}$ gives the reduced system for the multiplier update:
@@
The convergence of the Gauss-Seidel projection is of quasi-second order. Each update yields a monotone decrease with $O(1)$ work per constraint (hence $O(m)$ per sweep). With a strict tolerance and sufficiently small $\bm{\alpha}$, the final iterate satisfies the first-order optimality condition and realizes the hard-constrained projection. See Appendix E for details.

### v2

@@ Abstract
Biomolecular interaction modeling has been substantially advanced by foundation models, yet they often produce all-atom structures that violate basic steric feasibility. We address this limitation by enforcing physical validity as a strict constraint during both training and inference with a unified module. At its core is a differentiable projection that maps the provisional atom coordinates from the diffusion model to the nearest physically valid configuration. This projection is achieved using a Gauss-Seidel scheme, which exploits the locality and sparsity of the constraints to ensure stable and fast convergence at scale. By implicit differentiation to obtain gradients, our module integrates seamlessly into existing frameworks for end-to-end finetuning. With our Gauss-Seidel projection module in place, two denoising steps are sufficient to produce biomolecular complexes that are both physically valid and structurally accurate. Across six benchmarks, our $2$-step model achieves the same structural accuracy as state-of-the-art $200$-step diffusion baselines, delivering ${\sim}10\times$ wall-clock speedups while guaranteeing physical validity. The code is available at https://github.com/chensiyuan030105/ProteinGS.git.
@@ 1 Introduction
End-to-end, all-atom protein structure predictors that integrate deep learning with generative modeling are emerging as transformative tools for biomolecular interaction modeling (AlphaFold1; AlphaFold2; AlphaFold3; corso2023diffdock; Watson2023rfdiffusion). These systems achieve unprecedented accuracy in predicting arbitrary biomolecular complexes and are shaping the future of computational biology and drug discovery.
Contrary to their high structural accuracy, current predictors often fail to satisfy a conceptually simpler but basic requirement: the physical validity of the all-atom output. This failure, also commonly termed hallucination, appears as steric clashes, distorted covalent geometry, and stereochemical errors (Fig. 11). In practice, however, physical validity is a prerequisite: Atom-level physics violations hinder expert assessment (senior2020improved), undermine structure-based reasoning and experimental planning (lyu2019ultra), and also destabilize downstream computational analyses such as molecular dynamics (hollingsworth2018md; lindorff2011fastfolding).
The root cause of this problem lies in the design of current predictors: as generative models, they are trained to match the empirical distribution of known structures, without enforcing physical validity as a strict constraint in the training objective. Consequently, these models can assign non-zero probability to non-physical configurations, a flaw that persists even with large-scale training, both in data and model size, as observed in recent works (wohlwend2024boltz1; passaro2025boltz2; buttenschoen2024posebusters; bytedance2025protenix). Methods such as Boltz-1-Steering (wohlwend2024boltz1) reintroduce physics as inference-time guidance by biasing the sampling process towards physically valid regions. Such guidance can reduce violations, but still cannot guarantee validity: with finite guidance strength and update steps, invalid configurations remain reachable.
@@
Concretely, we instantiate this decoupling with a differentiable Gauss-Seidel projection placed after the denoiser. It solves a constrained optimization that projects all-atom coordinates to the nearest physically valid configuration. The constraints are tightly coupled through shared atoms and their number scales with the number of atoms, making the optimization problem large. Within the training loop, the module must also converge in a handful of iterations with low memory per forward and backward pass. These requirements make pure first-order gradient descent methods impractical, which require tiny steps and many iterations to converge. We therefore adopt a Gauss-Seidel scheme (saad2003iterative), which sweeps over all constraints for a few iterations, enforcing each locally by updating only the affected atoms. It leverages locality and sparsity of the constraints and thereby yields faster, more stable convergence than gradient descent. The projection module is differentiable via implicit differentiation. We integrate it as a drop-in layer within existing biomolecular interaction frameworks (e.g., Boltz (wohlwend2024boltz1)) and finetune end-to-end. At inference time, 2 denoising steps suffice to produce structurally accurate, physically valid protein complexes.
We evaluate our method on six protein-complex benchmarks: CASP15, Test (wohlwend2024boltz1), PoseBusters (buttenschoen2024posebusters), AF3-AB (AlphaFold3), dsDNA, and RNA-Protein (PXMeter). Comparisons include $200$-step generative models (Boltz-1, Boltz-1-Steering (wohlwend2024boltz1), Boltz-2 (passaro2025boltz2), and Protenix (bytedance2025protenix)) and the few-step baseline Protenix-Mini (gong2025protenixminiefficientstructurepredictor). Despite using only two denoising steps, our model achieves competitive structural accuracy while guaranteeing physical validity. For runtime, our model delivers a ${\sim}10\times$ wall-clock speedup over baselines. Our study largely closes the gap between guaranteed physical validity and state-of-the-art structural accuracy under few-step sampling, enabling 2-step all-atom predictions with an order-of-magnitude faster inference.
@@ 2 Related Work
Protein Structure Prediction. Deep learning-based foundation models are reshaping the task of protein structure prediction that produces 3D coordinates from input sequences (AlphaFold1; AlphaFold2). End-to-end variants extended this paradigm to complexes and large-scale training and inference (baek2021rosettafold; evans2022alphafoldmultimer). AlphaFold-3 further moved to all-atom biomolecular interaction modeling with a diffusion sampler (AlphaFold3), followed by improvements on generative models, such as Boltz-2 (passaro2025boltz2) and Protenix (bytedance2025protenix). These methods typically require hundreds of denoising steps. Few-step variants such as Protenix-Mini (gong2025protenixminiefficientstructurepredictor) reduce sampling to two steps while maintaining accuracy, yet still lack guarantees on physical validity. Boltz-1-Steering (wohlwend2024boltz1) is the first to explicitly target this issue by steering sampling with physics-informed potentials, but the guidance is soft and cannot preclude violations. By contrast, we introduce a Gauss-Seidel projection layer that enforces physical validity for both training and inference.
@@
Diffusion Models with Tilted Distribution Sampling. Diffusion models have demonstrated strong capabilities in producing samples given input conditions (Jonathan2022Denoising; Dhariwal2021Diffusion; song2021score). However, in many cases, users require the generated samples to satisfy certain specific constraints. There are primarily two types of constraint-aware diffusion methods: (i) Feynman-Kac (FK) steering-based conditional sampling methods (trippe2023diffusion; singhal2025general), which steer the sampling toward regions that meet desired conditions through reweighting of the path distribution in the diffusion process, and (ii) manifold-based diffusion models (debortoli2022riemannian; elhag2024manifold), which explicitly construct a constrained manifold, ensuring that generated samples inherently satisfy the desired constraints. However, FK-steering typically requires a large number of sampling steps, and explicit manifold construction can be computationally expensive when constraints are complex. We introduce a Gauss-Seidel projection module that enforces constraints in training and inference, thereby circumventing both excessive sampling iterations and the need for expensive manifold construction.
@@
Constraint Enforcement via Position-Based Dynamics. Enforcing physical constraints has long been central to physics simulation across computational physics and computer graphics. To handle boundary conditions (Macklin2014Unified), collisions, contact, friction (Baraff1998Robust), and bond stretching (Ryckaert1977Numerical), many methods have been developed to keep simulations physically consistent. Among these, position-based dynamics and its extensions (Muller2007PBD; Macklin2016XPBD; chen2024vertex) iteratively project particle positions to satisfy predefined constraints, proving effective for soft bodies, fluids, and coupled phenomena (Jan2014Survey). By updating positions directly instead of integrating stiff forces, these methods remain stable under large timesteps and avoid costly global solves. Since constraints are enforced through small, local projections, the overall complexity scales nearly linearly with the number of constraints, which aligns with our goal of generating proteins that respect local physical constraints.
Finetuning Diffusion Models. Finetuning has become essential for tailoring diffusion models to specific downstream tasks. To enhance controllability, many recent works employ reinforcement learning (RL) strategies: (i) methods such as DRaFT (clark2024directly) and DPPO (ren2024diffusion) treat the denoising trajectory as a policy to directly optimize reward functions; (ii) approaches like DPOK (fan2023dpok) and DiffPPO (Xiao2024DiffPPO) leverage preference data or human feedback. While effective at steering generation, these RL-based methods often incur high computational costs or exhibit sensitivity to the quality of the feedback signal. In contrast, we introduce a Gauss-Seidel projection during training to finetune the model via strictly defined physical constraints, circumventing the instability of policy gradient estimation.
@@ 3 Preliminary
All-Atom Diffusion Model. Proteins can be represented at all-atom resolution (AlphaFold3; wohlwend2024boltz1; passaro2025boltz2), where a structure is specified by the Cartesian coordinates of all atoms $\hat{\mathbf{x}}\in\mathbb{R}^{N\times 3}$, with $N$ the total number of atoms. This representation enables direct modeling of side-chain orientations, local packing, and explicit inter-atomic geometry. AlphaFold3 (AlphaFold3) introduced an atom-level diffusion approach that iteratively denoises atomic coordinates from Gaussian noise. Conditioned on features produced by the atom-attention encoder, MSA module, and PairFormer, a denoising network runs for hundreds of steps (typically 200) to generate the all-atom prediction $\hat{\mathbf{x}}$.
Physics-Guided Steering. All-atom diffusion sampling is stochastic and therefore cannot guarantee physical validity. To address this, Boltz-1-Steering (wohlwend2024boltz1), built upon the Feynman-Kac framework (singhal2025general), employs physics-informed potentials. At each denoising step, these potentials score the validity and tilt the learned denoising model toward a lower-energy configuration. The energy is a weighted sum of seven terms: tetrahedral atom chirality, bond stereochemistry, planarity of double bonds, internal ligand distance bounds, steric clashes avoidance, non-overlapping symmetric chains, and the preservation of covalently bonded chains. Since this tilted distribution cannot be sampled directly, the method uses importance sampling to draw multiple candidates, which are then refined with gradient descent updates to further reduce violations. Although this steering reduces physical errors, it remains a soft approach: it biases the sampling process toward validity during inference rather than embedding validity into the training.
@@ 4.1 Penalty-based Formulation for Physical Validity
where $\mathbf{C}(\cdot)\!:\!\mathbb{R}^{N\times 3}\rightarrow\mathbb{R}^{m}$ concatenates all constraints that characterize physical validity, and its zero set defines the feasible space. We follow Boltz-1-Steering (wohlwend2024boltz1) to define the constraints. Detailed formulation is provided in Appendix L. Since all terms are locally defined on pairs or four-atom groups, the total number of constraints $m$ scales with atoms and can be large. For instance, for PDB 8TGH which consists of $470$ residues with $7,082$ atoms, $m$ is ${\sim}13$M.
A standard approach to solving the constrained problem Eq. 1 is the penalty method (boyd2004convex), in which the constraints are incorporated as an exterior penalty into the objective:
@@
where we choose a quadratic penalty $E(\cdot)$ and $\bm{\alpha}\in\mathbb{R}^{m\times m}$ is a block-diagonal matrix of penalty coefficients. For sufficiently small $\bm{\alpha}$ (i.e., large weights $\bm{\alpha}^{-1}$), Eq. 2 serves as a numerical realization of the hard-constrained projection: the minimizer satisfies the constraint up to numerical tolerance. In our implementation, we set $\bm{\alpha}_{j}=10^{-7}$ or $10^{-6}$ to enforce a tight feasibility threshold.
@@ 4.2 Gauss-Seidel Constraint Projection
Solving Eq. 2 alone is classical and admits many off-the-shelf solvers. Embedding it as a differentiable module within a biomolecular modeling pipeline, however, imposes stricter requirements: very fast convergence per forward pass and training step, and stable convergence under stiff, tightly coupled constraints. Pure first-order gradient descent, as used in Boltz-1-Steering (wohlwend2024boltz1), is ill-suited: it requires small step sizes and prohibitively many iterations to converge to a zero constraint penalty, making unrolled end-to-end training slow. We therefore propose a Gauss-Seidel projection that exploits locality and sparsity by sequentially updating only the atoms affected by each constraint, thereby achieving fast and stable convergence.
@@
We then introduce the Lagrange multiplier $\bm{\lambda}(\mathbf{x})\!:=\!-\bm{\alpha}^{-1}\mathbf{C}(\mathbf{x})$ by following (stuart1998dynamical) and (servin2006interactive) to obtain the coupled system:
@@
with updates $\mathbf{x}^{(n+1)}=\mathbf{x}^{(n)}+\Delta\mathbf{x}$ and $\bm{\lambda}^{(n+1)}=\bm{\lambda}^{(n)}+\Delta\bm{\lambda}$. It can be shown that with initialization $\mathbf{x}^{(0)}=\hat{\mathbf{x}}$ and $\bm{\lambda}^{(0)}=\mathbf{0}$, the linearized iterates converge to a solution of the original nonlinear system Eq. 4. See Appendix I for a proof. Applying the Schur complement (fuzhen2005schur) with respect to $\mathbf{I}$ gives the reduced system for the multiplier update:
@@
The convergence of the Gauss-Seidel projection is of quasi-second order. Each update yields a monotone decrease with $O(1)$ work per constraint (hence $O(m)$ per sweep). With a strict tolerance and sufficiently small $\bm{\alpha}$, the final iterate satisfies the first-order optimality condition and realizes the hard-constrained projection. See Appendix J for details.